# The Amaterasu particle: constraining the superheavy dark matter origin of UHECRs

### v1

@@ Abstract
Amaterasu, the second most energetic ($244$ EeV) cosmic ray particle has been recently detected by the Telescope Array (TA) surface detector. The origin of the TA Amaterasu event is puzzling, as its arrival direction points back to a void in the local Universe, lacking conventional astrophysical ultra-high-energy (UHE) cosmic ray (CR) sources. Hence, we explore the possibility if this TA Amaterasu event could have originated from the decay of superheavy dark matter (SHDM) in the Milky Way. Such an origin also opens up multi-messenger detection channels in both UHE gamma-rays and UHE neutrinos. In this present work, using the TA Amaterasu event and the multi-messenger limits/sensitivities from various UHE telescopes, we place stringent constraints on the lifetime and mass of the SHDM. We find that the non-detection of the corresponding gamma-rays at the Pierre Auger Observatory (PAO) and the TA is in severe tension with the SHDM parameter space required to explain the TA Amaterasu event. Additionally, we extend the multi-messenger analysis to the future UHE gamma-ray and UHE neutrino telescopes such as PAO upgrade, GRAND 200k and IceCube-Gen2. We find that the bounds from the future neutrino telescopes will be able to compete with the present UHECR bounds. However, compared to the existing UHE gamma-ray bounds, the future PAO upgrade and the GRAND 200k gamma-ray detectors will improve the bounds on SHDM lifetime by at least one order of magnitude.
@@ I Introduction
The Telescope Array (TA) experiment has recently detected an ultra-high-energy cosmic ray (UHECR) event (named Amaterasu) of energy $2.44\times 10^{11}$ GeV which is the second highest energy UHECR ever detected [1]. The arrival direction (R.A.: $255.9\pm 0.6^{\circ}$, Dec.: $16.1\pm 0.5^{\circ}$) of this UHECR points back to a void in the local Universe rendering more questions than answers about its origin. As UHECRs above energy $5\times 10^{10}$ GeV propagating more than $\sim 10$ Mpc (mean free path) undergo energy losses due to the Greisen–Zatsepin–Kuzmin (GZK) effect [2, 3, 4, 5, 6, 7], the likely source of this event is expected to be located within a distance of $\sim 30$ Mpc [1]. The TA collaboration [1] analysed a catalogue of gamma-ray sources to identify the probable source and found that the known sources were [8, 9] located at distances beyond $30$ Mpc. One possibility could be that the primary UHECR is an iron nucleus at source, which might have undergone significant deflection in the extragalactic magnetic field (EGMF) during propagation to Earth [10, 11, 12]. Due to this, the actual source direction might be different than the arrival direction. Magnetic monopole has also been explored in the literature as a possible origin of the TA Amaterasu event [13, 14, 15, 16, 17, 18]. Another possible explanation of the origin could be attributed to decaying superheavy dark matter (SHDM) in the Milky Way or any local galaxy along the arrival direction. In this work, we examine if such a possibility of decaying SHDM in the Milky Way could give rise to the TA Amaterasu event.
@@
Contrary to the conventional production of ultra-high-energy cosmic rays through the diffusive shock acceleration (bottom-up), dark matter (DM) decay into standard model (SM) particles has also been hypothesised as a top-down UHECR production mechanism (See review[7]). In order to produce the observed energy of the TA Amaterasu event via DM decay, energy-momentum conservation states that the DM particles should have mass larger than $\sim 5\times 10^{11}$ GeV and can extend up to the energy scale ($10^{15}-10^{16}$ GeV), close to the typical grand unified theory (GUT) scale [19, 20, 21, 22, 23, 24]. Production of such SHDM may be possible via mechanisms such as vacuum or gravitational fluctuations in the early Universe during or after inflation [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 22, 36, 37, 38, 39]. Note that UHECRs can also be produced in principle from self-annihilation of SHDM particles. However, the resultant flux is density squared suppressed [40]. Our analysis is based on the Galactic SHDM decay possibility, the extra-galactic component of the flux is found to be negligible in comparison to the the dominant contribution from Galactic DM halo [41].
@@
SHDM particle ($\chi$) in the Milky Way halo can produce UHECR protons ($p$), ultra-high-energy (UHE) gamma-rays ($\gamma$) and UHE neutrinos ($\nu$) via the decay mode $\chi\to X\bar{X}$. Here, $X$ and $\bar{X}$ are the SM particle-antiparticle pairs [42]. For example, SHDM decaying via $b$ quark channel can produce these UHE secondaries: $\chi\to b\bar{b}\to p\bar{p}~{}\text{or}~{}\gamma\gamma~{}\text{or}~{}\nu\bar{\nu}$. The resulting abundances of gamma-rays and neutrinos generally dominate over that of protons for specific choice of the decay channel. This is in disagreement with the UHECR composition measurements [43, 44, 45], that hint towards dominance of nucleon over gamma-rays and neutrinos. Note that abundances of different nuclei from protons to iron in the UHECR spectrum is still an unresolved issue [43, 44, 45]. Because of the disagreement with observations, SHDM scenarios for UHECRs are tightly constrained up to the maximum observed CR energy, $\sim 10^{11}$ GeV. The conventional UHECR spectrum is heavily suppressed around $10^{11}$ GeV due to the GZK effect and falls with energy as $\sim E^{-5}$ [46]. Thus, the region above this energy becomes invaluable for probing scenarios of SHDM.
@@
In this work, we for the first time consider the origin of the recent TA Amaterasu event in light of SHDM decay. We put multi-messenger constraints on the lifetime and mass of the SHDM using UHECR, gamma-ray and neutrino telescopes. We also account for the low energy UHECR spectrum detected by the TA. We consider different decay modes for SHDM as well as a wide range ($10^{11}-10^{15}$ GeV) of SHDM mass. The constraints on the lifetime from gamma-ray non-detection limit of TA and PAO are found to be stronger than that of the UHECRs, discrediting the possibility of the TA Amaterasu event originating from SHDM decay. Furthermore, we also analyse the potential of the future telescopes mentioned above for probing SHDM lifetime through UHE gamma-ray and UHE neutrino detection. While future UHE neutrino detection are found to yield constraints comparable to that from UHECRs, UHE gamma-ray detection will lead to a few order of magnitudes improved constraints than that from UHECRs. Our constraints based on the TA Amaterasu event are consistent with those existing in the literature [41, 40, 59, 60, 61, 62, 63, 64, 65, 66, 50, 67, 68, 69, 70, 71, 72, 73, 74, 31, 73, 75, 76]. These works are based on various observations of UHECRs, UHE gamma-rays, and UHE neutrinos, but not the recent TA Amaterasu event.
The paper is organised as follows. In Sec. II, we describe the model adopted for flux calculation . In Sec. III, we present the results of our multi-messenger analysis for both present and future telescopes. Finally we summarize and discuss our main results in Sec. IV. The dependence of the results on various DM density profiles is discussed in Appendix A.
@@ II UHE flux from SHDM decay
Now, this diffuse flux of the secondary particles can be obtained for different decay modes of $\chi$, i.e., for different quarks, leptons and bosons for the $X$ and $\bar{X}$ combination (eg., $b\bar{b}$, $W^{+}W^{-}$, $\mu^{+}\mu^{-}$, $\nu\bar{\nu}$) in the decay process $\chi\to X\bar{X}$. In particular, the differential energy spectra ${dN_{i}(E)}/{dE}$ is estimated for these different decay modes.
@@
To demonstrate the flux estimation for the UHECRs, UHE gamma-rays and UHE neutrinos, we consider the example of $\chi\to b\bar{b}$ mode using Eq. (5) with $100\%$ branching fraction. The results are depicted in Fig. 1 together with the observational data and constraints of these UHE fluxes from the TA experiment. We consider two specific choices of $m_{\chi}$: $10^{12}$ GeV and $10^{14}$ GeV. The fluxes of all these UHE particles are inversely proportional (see Eq. (5) ) to $\tau_{\chi}$. Rather than fixing the parameter $\tau_{\chi}$ arbitrarily, we constrain the value of $\tau_{\chi}$ in a phenomenological manner. $\tau_{\chi}$ is estimated by normalising the SHDM derived flux of UHECRs, UHE gamma-rays and UHE neutrinos with the observed data or experimental upper limits on these fluxes. Since these fluxes as well as the observed flux/limits are all different from each other, they give rise to different values of $\tau_{\chi}$.
@@
The left plot of Fig. 1 shows the observed UHECR spectrum data by TA in black dots. The red data point represents the flux of the recently observed TA Amaterasu event of energy $2.44\times 10^{11}$ GeV [1]. The flux for this TA Amaterasu event is obtained by dividing the total number of events by the total exposure of the telescope [1]. The uncertainties in the flux correspond to the propagated uncertainties. Importantly, for the energies higher than $244$ EeV, we consider a conservative estimate that the maximum possible CR flux is not larger than the observed TA flux at the $244$ EeV event, shown by the purple dashed line. This is contrary to the harder CR spectra suggestions at the highest energies [43, 82]. For these maximum limits, we show the upper limit of $p+\bar{p}$ flux for two different DM masses ($m_{\chi}=10^{12}$ GeV and $m_{\chi}=10^{14}$ GeV) considering the decay of DM into $b\bar{b}$ channel. This establishes the lower bound on DM lifetime, as indicated by Eq. (5). For the DM masses of $10^{12}$ GeV and $10^{14}$ GeV, the observational constraints fix the $\tau_{\chi}$ at $5.5\times 10^{28}$ s and $7.5\times 10^{29}$ s, respectively. Note that the required energy for the TA Amaterasu event can not be created from SHDM with $m_{\chi}\lesssim 10^{12}$ GeV. This can be understood from the flux corresponding to $m_{\chi}=10^{12}$ GeV in the left plot of Fig. 1 as the TA Amaterasu event lies on the high energy tail of this flux. For $m_{\chi}\lesssim 10^{12}$ GeV, the flux will fall off at energies smaller than the energy of the TA Amaterasu event. Similarly, there exists an upper limit for the SHDM mass. For example, the flux corresponding to $m_{\chi}=10^{14}$ GeV is falling below the TA flux for our assumption of constant UHECR limit beyond the energy of the TA Amaterasu event. In particular, we find an upper limit on $m_{\chi}$ for this specific channel to be $m_{\chi}\sim 4\times 10^{13}$ GeV.
@@
In all the three plots of Fig. 1, we do not take into account the contribution from the conventional astrophysical as well as extragalactic SHDM sources. For UHECRs, the contribution of the extragalactic SHDM component is found to be negligible [40, 41]. Whereas, the contribution of the conventional astrophysical component might be significant in the UHECR spectral energy range below the energy of the TA Amaterasu event, i.e., $\lesssim 10^{11}$  GeV [41]. However, considering a dominant fraction of UHECRs from the conventional astrophysical origin allows only a little fraction of UHECRs from SHDM. This results in stringent constraints on the production of UHECRs via SHDM decay in the DM mass range $\lesssim 10^{12}$ GeV and do not yield a conservative description of the constraints on SHDM lifetime. In order to avoid this, we consider that the UHECRs in this energy range are solely due to SHDM decay. This assumption gives rise to a lower bound on the SHDM lifetime. For UHE gamma-rays, the contribution of the extragalactic component is negligible in the energy range of our interest in comparison to the Galactic SHDM component. This is because UHE gamma-rays from extragalactic sources are attenuated severely due to pair production losses in the cosmic microwave background (CMB) and the extragalactic radio background during propagation to Earth [4, 41]. For UHE neutrinos, the contribution from extragalactic sources can be important as unlike gamma-rays, neutrinos do not suffer any propagation losses. Inclusion of extragalactic sources will result in lowering of the UHE neutrino flux from SHDM decay in the Milky Way which, in turn, will give rise to larger SHDM lifetime. In order to be consistent with the assumption for UHECRs above, we do not include the extragalactic neutrino flux in our analysis. Due to these assumptions, the constraints obtained on the SHDM lifetime from different messengers can be comprehended as the most conservative constraints. In the following, we discuss these constraints for present and future telescopes in detail and expand the description to different decay channels.
@@ III Multi-messenger constraints on SHDM lifetime
UHECR telescopes and measurements: UHECRs were discovered by the Akeno Giant Air Shower Array (AGASA) [85] and the High Resolution Fly’s Eye (HiRes) [86, 87] telescopes about a couple of decades ago. In fact, HiRes had detected the most energetic ($320$ EeV) UHECRs till date [88]. Recent measurements of the UHECR spectrum above $0.1$ EeV by the PAO [47] and TA [89] experiments agree well with these old measurements except for some disagreements regarding mass composition and spectral index at extreme energies. In this work, we focus on the UHECR spectrum measured by TA and also on the recent TA Amaterasu event. In future, the Giant Radio Array for Neutrino Detection (GRAND) [57] will complement these existing experiments and may provide us with useful insights into UHECR source properties.

### v2

@@ Abstract
Amaterasu, the second most energetic ($244$ EeV) cosmic ray particle has been recently detected by the Telescope Array (TA) surface detector. The origin of the TA Amaterasu event is puzzling, as its arrival direction points back to a void in the local Universe, lacking conventional astrophysical ultra-high-energy (UHE) cosmic ray sources. Hence, we explore the possibility if this TA Amaterasu event could have originated from the decay of superheavy dark matter (SHDM) in the Milky Way. Such an origin also opens up multi-messenger detection channels in both UHE gamma-rays and UHE neutrinos. In this present work, using the TA Amaterasu event and the multi-messenger limits/sensitivities from various UHE telescopes, we place stringent constraints on the lifetime and mass of the SHDM. We find that the non-detection of the corresponding gamma-rays at the Pierre Auger Observatory (PAO) and the TA is in severe tension with the SHDM parameter space required to explain the TA Amaterasu event. Additionally, we extend the multi-messenger analysis to the future UHE gamma-ray and UHE neutrino telescopes such as PAO upgrade, GRAND 200k and IceCube-Gen2. We find that the bounds from the future neutrino telescopes will be able to compete with the present UHECR bounds. However, compared to the existing UHE gamma-ray bounds, the future PAO upgrade and the GRAND 200k gamma-ray detectors will improve the bounds on SHDM lifetime by at least one order of magnitude.
@@ I Introduction
The Telescope Array (TA) experiment has recently detected an ultra-high-energy cosmic ray (UHECR) event (named Amaterasu) of energy $2.44\times 10^{11}$ GeV which is the second highest energy UHECR ever detected [1]. The arrival direction (R.A.: $255.9\pm 0.6^{\circ}$, Dec.: $16.1\pm 0.5^{\circ}$) of this UHECR points back to a void in the local Universe rendering more questions than answers about its origin. As UHECRs above energy $5\times 10^{10}$ GeV propagating more than $\sim 10$ Mpc (mean free path) undergo energy losses due to the Greisen–Zatsepin–Kuzmin (GZK) effect [2, 3, 4, 5, 6, 7, 8, 9], the likely source of this event is expected to be located within a distance of $\sim 30$ Mpc [1]. The TA collaboration [1] analysed a catalogue of gamma-ray sources to identify the probable source and found that the known sources were [10, 11] located at distances beyond $30$ Mpc. One possibility could be that the primary UHECR is an iron nucleus at source, which might have undergone significant deflection in the extragalactic magnetic field (EGMF) during propagation to Earth [12, 13, 14]. Due to this, the actual source direction might be different than the arrival direction. Magnetic monopole has also been explored in the literature as a possible origin of the TA Amaterasu event [15, 16, 17, 18, 19, 20]. Another possible explanation of the origin could be attributed to decaying superheavy dark matter (SHDM) in the Milky Way or any local galaxy along the arrival direction. In this work, we examine if such a possibility of decaying SHDM in the Milky Way could give rise to the TA Amaterasu event.
@@
Contrary to the conventional production of UHECRs through the diffusive shock acceleration (bottom-up), dark matter (DM) decay into standard model (SM) particles has also been hypothesised as a top-down UHECR production mechanism (See review[7]). In order to produce the observed energy of the TA Amaterasu event via DM decay, energy-momentum conservation states that the DM particles should have mass larger than $\sim 5\times 10^{11}$ GeV and can extend up to the energy scale ($10^{15}-10^{16}$ GeV), close to the typical grand unified theory (GUT) scale [21, 22, 23, 24, 25, 26, 27, 28]. Production of such SHDM may be possible via mechanisms such as vacuum or gravitational fluctuations in the early Universe during or after inflation [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 26, 40, 41, 42, 43]. Note that UHECRs can also be produced in principle from self-annihilation of SHDM particles. However, the resultant flux is density squared suppressed [44]. Our analysis is based on the Galactic SHDM decay possibility, the extra-galactic component of the flux is found to be negligible in comparison to the the dominant contribution from Galactic DM halo [45].
@@
SHDM particle ($\chi$) in the Milky Way halo can produce UHECR protons ($p$), ultra-high-energy (UHE) gamma-rays ($\gamma$) and UHE neutrinos ($\nu$) via the decay mode $\chi\to X\bar{X}$. Here, $X$ and $\bar{X}$ are the SM particle-antiparticle pairs [46, 47, 48, 49, 50]. For example, SHDM decaying via $b$ quark channel can produce these UHE secondaries: $\chi\to b\bar{b}\to p\bar{p}~{}\text{or}~{}\gamma\gamma~{}\text{or}~{}\nu\bar{\nu}$. The resulting abundances of gamma-rays and neutrinos generally dominate over that of protons for specific choice of the decay channel. This is in disagreement with the UHECR composition measurements [51, 52, 53], that hint towards dominance of nucleon over gamma-rays and neutrinos. Note that abundances of different nuclei from protons to iron in the UHECR spectrum is still an unresolved issue [51, 52, 53]. Because of the disagreement with observations, SHDM scenarios for UHECRs are tightly constrained up to the maximum observed CR energy, $\sim 10^{11}$ GeV. The conventional UHECR spectrum is heavily suppressed around $10^{11}$ GeV due to the GZK effect and falls with energy as $\sim E^{-5}$ [54]. Thus, the region above this energy becomes invaluable for probing scenarios of SHDM.
@@
In this work, we for the first time consider the origin of the recent TA Amaterasu event in light of SHDM decay. We put multi-messenger constraints on the lifetime and mass of the SHDM using UHECR, gamma-ray and neutrino telescopes. We also account for the low energy UHECR spectrum detected by the TA. We consider different decay modes for SHDM as well as a wide range ($10^{11}-10^{15}$ GeV) of SHDM mass. The constraints on the lifetime from gamma-ray non-detection limit of TA and PAO are found to be stronger than that of the UHECRs, discrediting the possibility of the TA Amaterasu event originating from SHDM decay. Furthermore, we also analyse the potential of the future telescopes mentioned above for probing SHDM lifetime through UHE gamma-ray and UHE neutrino detection. While future UHE neutrino detection are found to yield constraints comparable to that from UHECRs, UHE gamma-ray detection will lead to a few order of magnitudes improved constraints than that from UHECRs. Our constraints based on the TA Amaterasu event are consistent with those existing in the literature [45, 44, 67, 68, 69, 70, 71, 72, 73, 74, 58, 75, 76, 77, 78, 79, 80, 81, 82, 35, 81, 83, 84, 85]. These works are based on various observations of UHECRs, UHE gamma-rays, and UHE neutrinos, but not the recent TA Amaterasu event.
The paper is organised as follows. In Sec. II, we describe the model adopted for flux calculation . In Sec. III, we present the results of our multi-messenger analysis for both present and future telescopes. Finally we summarize and discuss our main results in Sec. IV. The dependence of the results on various DM density profiles is discussed in Appendix A and the constraints from the Galactic center are discussed in Appendix B.
@@ II UHE flux from SHDM decay
Now, this diffuse flux of the secondary particles can be obtained for different decay modes of $\chi$, i.e., for different quarks, leptons and bosons for the $X$ and $\bar{X}$ combination (eg., $b\bar{b}$, $W^{+}W^{-}$, $\mu^{+}\mu^{-}$, $\nu\bar{\nu}$) in the decay process $\chi\to X\bar{X}$. In particular, the differential energy spectra ${dN_{i}(E)}/{dE}$ is estimated for these different decay modes. To demonstrate the flux estimation for the UHECRs, UHE gamma-rays and UHE neutrinos, we consider the example of $\chi\to b\bar{b}$ mode using Eq. (5) with $100\%$ branching fraction. The results are depicted in Fig. 1 together with the observational data and constraints of these UHE fluxes from the TA experiment. We consider two specific choices of $m_{\chi}$: $10^{12}$ GeV and $10^{14}$ GeV. The fluxes of all these UHE particles are inversely proportional (see Eq. (5) ) to $\tau_{\chi}$. Rather than fixing the parameter $\tau_{\chi}$ arbitrarily, we constrain the value of $\tau_{\chi}$ in a phenomenological manner. $\tau_{\chi}$ is estimated by normalising the SHDM derived flux of UHECRs, UHE gamma-rays and UHE neutrinos with the observed data or experimental upper limits on these fluxes. Since these fluxes as well as the observed flux/limits are all different from each other, they give rise to different values of $\tau_{\chi}$.
@@
The left plot of Fig. 1 shows the observed UHECR spectrum data by TA in black dots. The red data point represents the flux of the recently observed TA Amaterasu event of energy $2.44\times 10^{11}$ GeV [1]. The flux for this TA Amaterasu event is obtained by dividing the total number of events by the total exposure of the telescope [1]. The uncertainties in the flux correspond to the propagated uncertainties. Importantly, for the energies higher than $244$ EeV, we consider a conservative estimate that the maximum possible CR flux is not larger than the observed TA flux at the $244$ EeV event, shown by the purple dashed line. This is contrary to the harder CR spectra suggestions at the highest energies [51, 91]. For these maximum limits, we show the upper limit of $p+\bar{p}$ flux for two different DM masses ($m_{\chi}=10^{12}$ GeV and $m_{\chi}=10^{14}$ GeV) considering the decay of DM into $b\bar{b}$ channel. This establishes the lower bound on DM lifetime, as indicated by Eq. (5). For the DM masses of $10^{12}$ GeV and $10^{14}$ GeV, the observational constraints fix the $\tau_{\chi}$ at $5.5\times 10^{28}$ s and $7.5\times 10^{29}$ s, respectively. Note that the required energy for the TA Amaterasu event can not be created from SHDM with $m_{\chi}\lesssim 10^{12}$ GeV. This can be understood from the flux corresponding to $m_{\chi}=10^{12}$ GeV in the left plot of Fig. 1 as the TA Amaterasu event lies on the high energy tail of this flux. For $m_{\chi}\lesssim 10^{12}$ GeV, the flux will fall off at energies smaller than the energy of the TA Amaterasu event. Similarly, there exists an upper limit for the SHDM mass. For example, the flux corresponding to $m_{\chi}=10^{14}$ GeV is falling below the TA flux for our assumption of constant UHECR limit beyond the energy of the TA Amaterasu event. In particular, we find an upper limit on $m_{\chi}$ for this specific channel to be $m_{\chi}\sim 4\times 10^{13}$ GeV. However, later in Sec. III, we show that this range of $m_{\chi}$ is in significant conflict with the UHE gamma-ray constraints from different telescopes.
@@
In all the three plots of Fig. 1, we do not take into account the contribution from the conventional astrophysical as well as extragalactic SHDM sources. For UHECRs, the contribution of the extragalactic SHDM component is found to be negligible [44, 45]. Whereas, the contribution of the conventional astrophysical component might be significant in the UHECR spectral energy range below the energy of the TA Amaterasu event, i.e., $\lesssim 10^{11}$  GeV [45]. However, considering a dominant fraction of UHECRs from the conventional astrophysical origin allows only a little fraction of UHECRs from SHDM. This results in stringent constraints on the production of UHECRs via SHDM decay in the DM mass range $\lesssim 10^{12}$ GeV and do not yield a conservative description of the constraints on SHDM lifetime. In order to avoid this, we consider that the UHECRs in this energy range are solely due to SHDM decay. This assumption gives rise to a lower bound on the SHDM lifetime. For UHE gamma-rays, the contribution of the extragalactic component is negligible in the energy range of our interest in comparison to the Galactic SHDM component. This is because UHE gamma-rays from extragalactic sources are attenuated severely due to pair production losses in the cosmic microwave background (CMB) and the extragalactic radio background during propagation to Earth [4, 45]. Note that, the attenuation of Galactic UHE gamma-rays due to the pair production losses is negligible [45, 85]. For UHE neutrinos, the contribution from extragalactic sources can be important as unlike gamma-rays, neutrinos do not suffer any propagation losses. Inclusion of extragalactic sources will result in lowering of the UHE neutrino flux from SHDM decay in the Milky Way which, in turn, will give rise to larger SHDM lifetime. In order to be consistent with the assumption for UHECRs above, we do not include the extragalactic neutrino flux in our analysis. Due to these assumptions, the constraints obtained on the SHDM lifetime from different messengers can be comprehended as the most conservative constraints. In the following, we discuss these constraints for present and future telescopes in detail and expand the description to different decay channels.
@@ III Multi-messenger constraints on SHDM lifetime
UHECR telescopes and measurements: UHECRs were discovered by the Akeno Giant Air Shower Array (AGASA) [94] and the High Resolution Fly’s Eye (HiRes) [95, 96] telescopes about a couple of decades ago. In fact, HiRes had detected the most energetic ($320$ EeV) UHECRs till date [97]. Recent measurements of the UHECR spectrum above $0.1$ EeV by the PAO [55] and TA [98] experiments agree well with these old measurements except for some specific disagreements regarding mass composition and spectral index at extreme energies. In this work, we focus on the UHECR spectrum measured by TA and also on the recent TA Amaterasu event. In future, the Giant Radio Array for Neutrino Detection (GRAND) [65] will complement these existing experiments and may provide us with useful insights into UHECR source properties.